# System performance of a cryogenic test-bed for the time-division multiplexing readout for NewAthena X-IFU

### v2

@@ Abstract
The X-ray Integral Field Unit (X-IFU) is an instrument of ESA’s future NewAthena space observatory, with the goal to provide high-energy resolution ( 4 eV at X-ray energies up to 7 keV) and high-spatial resolution (9”) spectroscopic imaging over the X-ray energy range from 200 eV to 12 keV, by means of an array of about 1500 transition-edge sensors (TES) read out via SQUID time-division multiplexing (TDM). In 2022, to aid in the transfer of TDM readout technology from the laboratory toward flight hardware, our team commissioned a new TDM-based laboratory test-bed at SRON. This setup hosts an array of m2 TESs that are read out via 2-column 32-row TDM. A system component that is critical to high-performance operation is the wiring harness that connects the room-temperature electronics to the cryogenic readout componentry. In November 2023, we implemented a re-designed flex harness, which in the SRON test-bed has a length close to what is envisioned for the X-IFU flight harness. We report here on our characterization of the TDM system with the new flex harness, which allowed the system to achieve a co-added energy resolution at a level of 2.7 eV FWHM at 6 keV via 32-row readout. In addition, we provide an outlook on the upcoming integration of TDM readout into the X-IFU Focal-Plane Assembly Development Model.
@@
Time-division multiplexing, transition-edge sensor, X-IFU, X-ray spectroscopy1 Introduction
@@
The X-IFU will make use of 1500 transition-edge sensors (TES) micro-calorimeters [3], organized in an hexagonal array covering a field of view of 4’ with a pixel size of m corresponding to 5” PSF angular resolution. The sensors will be cooled down to a base temperature of 50 mK by means of a multi-stage Adiabatic Demagnetization Refrigerator (ADR) and are designed to operate in the soft X-ray energy band, between 200 eV and 12 keV, with a predicted energy resolution at instrument level better than 4 eV for energies up to 7 keV.
@@
In TDM readout, the TES are continuously dc-biased and are each coupled to a first stage SQUID (SQ1), activated one at a time via a flux-actuated switch (FAS). SQ1 signals are amplified by a SQUID series array (SSA) and digitized by room temperature electronics. Each SQ1 represents a “row” and the FAS are activated by sequential square wave signals, or row-addressing signals (RAS), generated by the digital readout electronics (DRE). The SQ1s are hosted on “MUX” chips, together with the shunt resistance to voltage bias the TES and Nyquist inductors to limit the pixel bandwidth. Each row is active for a certain “row time”, all the rows are read out within one “frame” and the columns are read out in parallel. For X-IFU, a 48-row 32-column TDM readout is envisaged, with row time of 160 ns.
@@
SRON is responsible for the design, development and delivery of X-IFU’s Focal Plane Assembly (FPA). The first iteration of the FPA Development Model (FPA-DM) was employing a TES array[6] and a frequency-domain multiplexing (FDM)[7] technology, both developed at SRON. It was successfully used to validate its thermo-mechanical and magnetic environment. A second iteration of the FPA-DM, called DM1.1[8], is foreseen before upscaling to the FPA Engineering Model, which will employ a TDM architecture. For the assembly and test campaign of the DM1.1, expected to start in Q1 2025, SRON will receive key components from several external partners: the TES array and the digital room temperature electronics (DRE) from NASA-GSFC (USA), the MUX chips from NIST (USA) and the warm front-end electronics (WFEE) from APC-CNRS (France).
@@
In late 2022, our team assembled components from SRON, NIST and NASA-GSFC into a new cryogenic test-bed at SRON, as described in Section 2. The motivation was to transfer TDM technology from the development laboratory to a more flight-like configuration ahead of its implementation in the DM1.1. Driven by the NewAthena spacecraft geometry, the flight-like TDM system will require a significantly longer wiring harness that links the room-temperature electronics to the cryogenic componentry that is typically used in laboratory TDM systems. Thus, this new test-bed at SRON uses a longer wiring harness. An upgrade in late 2023 introduced a re-designed version of the long harness (see Section 2.1). Our main recent goal was to reproduce in this new system the nominal spectral performance of the previous laboratory systems: 2.8 eV FWHM or better energy resolution at 6 keV under 2-column 32-row multiplexing.
@@
In this paper we report on the experimental activities performed with such cryogenic test-bed at SRON. In Section 2 we describe both the hardware and software composing the experimental setup. In Section 3 we describe the characterization tests performed to optimize and debug the system. In Section 4 we show the performance demonstration, in agreement with the nominal specifications. Finally, in Section 5 we summarize our results and give a prospect for future activities.
@@ 2 Description of the setup
For ease of reading, we subdivide the components of the test-bed into three categories: cryogenic assembly, room temperature and control software.
@@ 2.1 Cryogenic assembly
A “50 mK snout” [9] hosts the TES [10] and MUX chips [11] ([A] in Figure 1). The detector array, provided by NASA-GSFC, is a kilo-pixel chip with TES made of m2 Mo/Au bilayer and Bi/Au absorber, with m and mK. The MUX chips, manufactured at NIST, are of design “mux20a”, in which each SQ1 is a two-SQUID series array. Accompanying “interface” chips, which contain the circuit elements for the TES-bias loop, are designed with nH and .
@@
The snout is suspended from the mixing chamber of the refrigerator using a two-stage Kevlar vibration isolation system [13] ([D] in Figure 1) to prevent power dissipation in the TES due to mechanical contributions in the system and mitigate low-frequency noise in the detectors’ noise from such sources, like the Pulse Tube. Thermal anchoring to the mixing chamber is performed via oxygen-free high-conductivity copper braids. Temperature is monitored via a ruthenium oxide thermometer and controlled via a PID (Proportional-Integral-Derivative) loop through a 750 heater installed on the snout’s mechanical support. An operating temperature of 50 mK is set and typical RMS temperature stability achieved is at the level of 1 K.
@@ 3 K board and inter-stage harness
The “3 K board” ([E] in Figure 1) hosts the SA20a series SQUID array (SSA), the second-stage amplifiers for the TES signals. Both the SSA and board assembly have been manufactured at NIST. The SSA chip is mounted within casing for magnetic shielding ([F] in Figure 1), made of an external layer of niobium and an internal layer of cryogenic mu-metal [12]. The purpose of this board is as well to provide signal routing and filtering. A number of utility connections are also available, for control of the superconducting magnet coil.
@@
This structure is motivated by necessity to limit the physical distance between the SSA and the MUX chips, to allow for sufficient interstage bandwidth between the two SQUID amplification stages, with the cut-off frequency given by , with the input inductance of the SSA, the dynamic output resistance of the SQ1 (or MUX SQUID) and the inductive contributions coming from the 50 mK to 3 K harness, PCB striplines, bonding wires, etc. Given the general requirement MHz (for a row time of 160 ns) and typical values of nH and 15 , practical implementations require such harness length to be at a level of 30 cm. In this test-bed, such harness is constituted by niobium-titanium twisted pair looms, with a thermal break in between via inter-stage filter cards ([G] in Figure 1) thermally anchored to the heat exchanger plate ( 50 mK).
@@
The connection between the “3 K board” and the warm analog electronics (“Tower”, see Subsection 2.2) is performed via flex harness ([H] in Figure 1). There are a total of 10 flexes, each constituted by eight supply and return conductor pairs, as diagrammed in Figure 2. The flexes are stacked and thermalised via dedicated heatsink plates at 3 K and 50 K. The interface of the flexes with the Tower electronics is described in Subsection 2.2.
@@
For both designs, the flex length is at a level of 1.5 m, significantly longer than the standard 40 cm used in ground-based TDM systems. The motivation for this difference was the reproduction of the nominal performance of a TDM system in a SRON dilution refrigerator, where the distance between the 3 K stage and room-temperature is larger than in the adiabatic demagnetization refrigerators (ADR) available at NIST and NASA-GSFC. Additionally, this is seen as an intermediate characterization of the system with a harness length closer to the 2.3 m required for X-IFU.
The dimensioning of the signal and return lines for the old flex design (Figure 2a) was chosen to keep the electrical resistance below 20 to be compliant with the requirements for the electronics, as well as limiting the capacitance between signal and return lines to prevent bandwidth loss. This however resulted in a larger cable inductance which enhanced electrical crosstalk in the system, inducing large electrical oscillations (“ringing”) with large amplitude, ultimately affecting the system performance already in first validation tests at NIST.
This led to the new design of the signal and return lines (Figure 2b), similar to the standard dimensioning for the 40 cm length. Warm bench test measurements done at NIST showed that the level of electrical crosstalk in the harness was reduced by a factor of 20, which significantly reduced the amplitude of the ringing observed in the old flex. The wider signal track resulted however in a larger capacitance, which caused a reduction in harness bandwidth. The consequence of this is discussed in Section 3.
@@ 2.2 Room temperature electronics
The control of the bias boards is performed via a serial protocol, executed from the measurement computer via an optical fiber connection, whereas the connections from the Tower to the DRE is performed via high-bandwidth HDMI cables for the RAS and BNC cables for the ADC and feedback signals.
@@
The DRE [16] is developed at NASA-GSFC and is composed of two elements: the Column Box and the Row Box ([J] in Figure 1). The Row Box generates the RAS to subsequently activate the rows of each column. Synchronicity between each row’s activation window (row time) and ADC sampling performed by the Column box is ensured via several clocks (frame clock, line clock) generated from a master clock at 245.76 MHz. The Column box furthermore handles the digital feedback for the FLL as well as DAQ back-end for data streaming.
@@ 2.3 Control software
Both the DRE from NASA-GSFC and the Tower from NIST have their own custom software package for measurement and control. The Tower software is used to provide bias to the SSA and SQ1 SQUIDs and to the TES arrays, while the DRE software handles all the TDM control aspects, including row-timing and FLL settings, as well as measurement settings.
@@
This “TES-TDM Readout Software” is built on SRON’s generic ground support software platform, Egress. It is designed to be scalable to a high number of readout channels, with components that require high performance implemented in C++. In addition to automated measurement scripts, a text-based user interface is provided.
@@ 3.1 EMI optimization
To prevent ground loops, a so-called “star” grounding scheme is used, where the top of the cryostat is used as ground reference. All measurement electronics are grounded via thick copper braids to such reference. Electronics modules, including the Column and Row Boxes, are powered on via an isolation transformer to decouple the ground reference of the TDM to the ground of the wall socket. Dedicated wall sockets in the laboratory, connected to separated electrical units in the building, are used for powering on the readout electronics of cryogenic setups. The top of the cryostat is then connected to a “clean” ground reference in the laboratory via a thick copper grounding cable. Furthermore, to electrically decouple the connection between the Tower and the Column box to the measurement computer, optical fibers are used.
Despite these precautions, strong noise lines at 50 Hz and higher harmonics have been initially observed in SQUID noise measurements. This was identified as due to electromagnetic interference (EMI) in the loop constituted by the 3 m HDMI and BNC cables connecting the DRE to the Tower. In fact, it was observed that bundling and intertwining the cables together, to minimize the area of the loop and reject common-mode noise, significantly reduced the amplitude of such noise lines. Finally, enclosing the cable bundle in aluminium foil almost completely removed the 50 Hz noise lines. High-frequency (several kHz) lines, thought to be spurs from the DRE, remained present in the system. However, their frequency is sufficiently higher with respect to the 250 Hz of the detectors electrical bandwidth at the quiescence operating point to consider their impact negligible.
@@ 3.2 Magnetic field offset
As it is customary for TES X-ray micro-calorimeter setups, to achieve the optimal spectral performance from the system, the behavior of the detectors under external magnetic field has to be characterized. This is because the TES critical current is magnetic field dependent, as the system formed by the TES bilayer and the niobium leads constitutes an SS’S Josephson junction. It is typically observed that the best operation is obtained when no magnetic fields perpendicular to the bilayer are present.
These fields might be caused by external sources such as pulse tubes, Earth magnetic field, EMI from laboratory equipment, etc. Furthermore, TES under dc-bias suffer from the effects of self-field, the magnetic field generated by the loop made by the leads and the TES itself [17]. Such magnetic field adds to the residual field at the TES array and is bias point dependent. The external magnetic fields are typically compensated with a superconducting coil mounted just below the TES array, as also envisaged for X-IFU.
@@
We found the average of the cancelling fields measured with both methods to be consistent. We measured an average cancelling magnetic field at a level of 150 nT with no evident gradient across the array, with few TES showing a variation in residual field with respect to neighbouring pixels (see Figure 4). The reason for this variation is not identified, but it could be due to either uncertainty in the measurement or a real effect in the TES array due to non-perfect uniformity. For reference, this value is nevertheless within the T requirement for X-IFU for magnetic-field gradient across the TES array. Furthermore, this value of average residual field is comparable with that typically measured for other TES setups readout under Frequency Domain Multiplexing hosted in the same cryostat [18]. The minimization of magnetic field gradients is important for flight operation, because TES pixels within a TDM readout column will share their dc-voltage bias and thus cannot be tuned individually to compensate for eventual different magnetic-field offsets.

### v3

@@ Abstract
The X-ray Integral Field Unit (X-IFU) is an instrument of ESA’s future NewAthena space observatory, with the goal to provide high-energy resolution ( 4 eV at X-ray energies up to 7 keV) and high-spatial resolution (9”) spectroscopic imaging over the X-ray energy range from 200 eV to 12 keV, by means of an array of about 1500 transition-edge sensors (TES) read out via SQUID time-division multiplexing (TDM). A TDM-based laboratory test-bed has been assembled at SRON, hosting an array of m2 TESs that are read out via 2-column 32-row TDM. A system component that is critical to high-performance operation is the wiring harness that connects the room-temperature electronics to the cryogenic readout componentry. We report here on our characterization of such a test-bed, whose harness has a length close to what envisioned for X-IFU, which allowed to achieve a co-added energy resolution at a level of 2.7 eV FWHM at 6 keV via 32-row readout. In addition, we provide an outlook on the integration of TDM readout into the X-IFU Focal-Plane Assembly Development Model.
@@
Time-division multiplexing, transition-edge sensor, X-IFU, X-ray spectroscopyThis paper has been accepted for publication in Journal of Astronomical Telescopes, Instrumentation and Telescopes.
1 Introduction
@@
The X-IFU will make use of 1500 transition edge-sensors (TES) micro-calorimeters [3], organized in a hexagonal array covering a field of view of 4’ with a pixel size of m corresponding to 5” PSF angular resolution. The sensors will be cooled down to a base temperature of 50 mK by means of a multi-stage Adiabatic Demagnetization Refrigerator (ADR) and are designed to operate in the soft X-ray energy band, between 200 eV and 12 keV, with a predicted energy resolution at instrument level better than 4 eV for energies up to 7 keV.
@@
In TDM readout, the TESs are continuously dc-biased and are each coupled to a first stage SQUID (SQ1), activated one at a time via a flux-actuated switch (FAS). SQ1 signals are amplified by a SQUID series array (SSA) and digitized by room temperature electronics. Each SQ1 represents a “row” and the FASs are activated by sequential square wave signals, or row-addressing signals (RAS), generated by the digital readout electronics (DRE). The SQ1s are hosted on “MUX” chips, together with the shunt resistance to voltage bias the TES and Nyquist inductors to limit the pixel bandwidth. Each row is active for a certain “row time”, all the rows are read out within one “frame” and the columns are read out in parallel. An explicative depiction of the operation of a TDM readout for micro-calorimeters can be found in Figure 8 of the recently published review “Signal Readout for Transition-Edge Sensor X-ray Imaging Spectrometers”[4]. For X-IFU, a 48-row 32-column TDM readout is envisaged, with row time of 160 ns.
@@
SRON is responsible for the design, development and delivery of X-IFU’s Focal Plane Assembly (FPA). The first iteration of the FPA Development Model (FPA-DM) was employing a TES array[6] and a frequency-domain multiplexing (FDM)[7] technology, both developed at SRON. It was successfully used to validate its thermo-mechanical and magnetic environment. A second iteration of the FPA-DM, called DM1.1[8], is foreseen before upscaling to the FPA Engineering Model, which will employ a TDM architecture. For the assembly and test campaign of the DM1.1, expected to start in Q2 2025, SRON will receive key components from several external partners: the TES array and the digital room temperature electronics (DRE) from NASA-GSFC (USA), the MUX chips from NIST (USA) and the warm front-end electronics (WFEE) from APC-CNRS (France).
@@
In late 2022, our team assembled components from SRON, NIST and NASA-GSFC into a new cryogenic test-bed at SRON, as described in Section 2. The motivation was to transfer TDM technology from the development laboratory to a more flight-like configuration ahead of its implementation in the DM1.1. Driven by the NewAthena spacecraft geometry, the flight-like TDM system will require a significantly longer wiring harness, linking the room-temperature electronics to the cryogenic componentry, than what is typically used in laboratory TDM systems.
Such a harness is critical for several aspects of the system operation, including crosstalk, low-frequency signal transfer from the SSA SQUID to the low-noise amplifier (LNA) at the room temperature electronics, the noise contribution of the harness resistance and the open-loop bandwidth of the signal transfer. Parameters such as harness length and termination on both the cold and the warm part of the readout have an impact on the aforementioned aspects in the system.
This new test-bed at SRON uses a longer wiring harness than laboratory TDM systems. An upgrade in late 2023 introduced a re-designed version of the long harness (see Section 2.1). Our main recent goal was to reproduce in this new system the nominal spectral performance of the previous laboratory systems: 2.8 eV FWHM or better energy resolution at 6 keV under 2-column 32-row multiplexing.
@@
In this paper we report on the experimental activities performed with such a cryogenic test-bed at SRON. In Section 2 we describe both the hardware and software composing the experimental setup. In Section 3 we describe the characterization tests performed to optimize and debug the system. In Section 4 we show the performance demonstration, in agreement with the nominal specifications. Finally, in Section 5 we summarize our results and give a prospect for future activities.
@@ 2 Description of the setup
For ease of reading, we subdivide the components of the test-bed into three categories: cryogenic assembly, room temperature electronics and control software.
@@ 2.1 Cryogenic assembly
A “50 mK snout” [9] hosts the TES [10] and MUX chips [11] ([A] in Figure 1). The detector array, provided by NASA-GSFC, is a kilo-pixel chip with each TES made of m2 Mo/Au bilayer and Bi/Au absorber, with m, pJ/K at 100 mK, pW/K at , mK. The MUX chips, manufactured at NIST, are of design “mux20a”[12], in which each SQ1 is a two-SQUID series array. Accompanying “interface” chips, which contain the circuit elements for the TES-bias loop, are designed with nH and .
@@
The snout is suspended from the mixing chamber of the refrigerator using a two-stage Kevlar vibration isolation system [14] ([D] in Figure 1) to prevent power dissipation in the TES array due to mechanical contributions in the system and mitigate low-frequency contributions to the detectors’ noise from such sources, like the Pulse Tube. Thermal anchoring to the mixing chamber is performed via oxygen-free high-conductivity copper braids. Temperature is monitored via a ruthenium oxide thermometer and controlled via a PID (Proportional-Integral-Derivative) loop through a 750 heater installed on the snout’s mechanical support. An operating temperature of 50 mK is set and the typical RMS temperature stability achieved is at the level of 1 K.
@@ 4 K board and inter-stage harness
The “4 K board” ([E] in Figure 1) hosts the SA20a series SQUID array (SSA), the second-stage amplifiers for the TES signals. Both the SSA and board assembly have been manufactured at NIST. The SSA chip is mounted within casing for magnetic shielding ([F] in Figure 1), made of an external layer of niobium and an internal layer of cryogenic mu-metal [13]. The purpose of this board is as well to provide signal routing and filtering. A number of utility connections are also available, for control of the superconducting magnet coil.
@@
This structure is motivated by necessity to limit the physical distance between the SSA and the MUX chips, to allow for sufficient interstage bandwidth between the two SQUID amplification stages, with the cut-off frequency given by , with the input inductance of the SSA, the dynamic output resistance of the SQ1 (or MUX SQUID) and the inductive contributions coming from the 50 mK to 3 K harness, PCB striplines, bonding wires, etc. Given the general requirement MHz (for a row time of 160 ns) and typical values of nH and 15 , practical implementations require such harness length to be at a level of 30 cm. In this test-bed, such a harness is constituted by niobium-titanium twisted pair looms, with a thermal break in between via inter-stage filter cards ([G] in Figure 1) thermally anchored to the heat exchanger plate ( 50 mK).
@@
The connection between the “4 K board” and the warm analog electronics (“Tower”, see Subsection 2.2) is performed via the flex harness ([H] in Figure 1). There are a total of 10 flexes, each constituted by eight supply and return conductor pairs, as diagrammed in Figure 2. The flexes are stacked and thermalised via dedicated heatsink plates at 3 K and 50 K. The interface of the flexes with the Tower electronics is described in Subsection 2.2.
@@
For both designs, the flex length is at a level of 1.5 m, significantly longer than the standard 40 cm used in ground-based TDM systems. The motivation for this difference was the reproduction of the nominal performance of a TDM system in an SRON dilution refrigerator, where the distance between the 3 K stage and room-temperature is larger than in the adiabatic demagnetization refrigerators (ADR) available at NIST and NASA-GSFC. Additionally, this is seen as an intermediate characterization of the system with a harness length closer to the 2.3 m required for X-IFU.
The old flex design employs symmetrical supply and return conductor pairs, constituted by 150 m Cu lines with 2.54 mm pair pitch. The dimensioning of the signal and return lines for the old flex design (Figure 2a) was chosen to keep the electrical resistance below 20 to be compliant with the requirements for the electronics, as well as limiting the capacitance between signal and return lines to prevent bandwidth loss. This however resulted in a larger cable inductance which enhanced electrical crosstalk in the system, inducing large electrical oscillations (“ringing”) with large amplitude, ultimately affecting the system performance already in first validation tests at NIST.
This led to the new design of the signal and return lines (Figure 2b), similar to the standard dimensioning for the 40 cm length. In this case, two CuNi 706 conductor tracks are asymmetrical, with a 460 m wide supply line and a 2 mm wide return line, still with a 2.54 mm pair pitch. Warm bench test measurements done at NIST showed that the level of electrical crosstalk in the harness was reduced by a factor of 20, which significantly reduced the amplitude of the ringing observed in the old flex. The wider signal track resulted however in a larger capacitance, which caused a reduction in harness bandwidth. The consequence of this is discussed in Section 3.
@@ 2.2 Room temperature electronics
The control of the bias boards is performed via a serial protocol, executed from the measurement computer via an optical fiber connection, whereas the connections from the Tower to the DRE are via high-bandwidth HDMI cables for the RAS and BNC cables for the ADC and feedback signals.
@@
The DRE [17] is developed at NASA-GSFC and is composed of two elements: the Column Box and the Row Box ([J] in Figure 1). The Row Box generates the RAS to subsequently activate the rows of each column. Synchronicity between each row’s activation window (row time) and ADC sampling performed by the Column box is ensured via several clocks (frame clock, line clock) generated from a master clock at 245.76 MHz. The Column box furthermore handles the digital feedback for the FLL as well as the DAQ back-end for data streaming.
@@ 2.3 Control software
Both the DRE from NASA-GSFC and the Tower from NIST have their own custom software package for measurement and control. The Tower software is used to provide bias to the SSA and SQ1 SQUIDs and to the TES array, while the DRE software handles all the TDM control aspects, including row-timing and FLL settings, as well as measurement settings.
@@
This “TES-TDM Readout Software” is built on SRON’s generic ground support software platform. It is designed to be scalable to a high number of readout channels, with components that require high performance implemented in C++. In addition to automated measurement scripts, a text-based user interface is provided.
@@ 3.1 EMI optimization
To prevent ground loops, a so-called “star” grounding scheme is used, where the top of the cryostat is used as ground reference. All measurement electronics are grounded via thick copper braids to such reference. Electronics modules, including the Column and Row Boxes, are powered on via an isolation transformer to decouple the ground reference of the TDM to the ground of the wall socket. Dedicated wall sockets in the laboratory, connected to separated electrical units in the building, are used for powering on the readout electronics of cryogenic setups. The top of the cryostat is then connected to a “clean” ground reference in the laboratory via a thick copper grounding cable. Furthermore, to electrically decouple the connections from the Tower and the Column box to the measurement computer, optical fibers are used.
Despite these precautions, strong noise lines at 50 Hz and higher harmonics have been initially observed in SQUID noise measurements. This was identified as due to electromagnetic interference (EMI) in the loop constituted by the 3 m HDMI and BNC cables connecting the DRE to the Tower. In fact, it was observed that bundling and intertwining the cables together, to minimize the area of the loop and reject common-mode noise, significantly reduced the amplitude of such noise lines. Finally, enclosing the cable bundle in aluminium foil almost completely removed the 50 Hz noise lines. High-frequency (several kHz) lines, thought to be spurs from the DRE, remained present in the system. However, their frequency is sufficiently higher with respect to the 250 Hz of the detector’s electrical bandwidth at the quiescence operating point to consider their impact negligible.
@@ 3.2 Magnetic field offset
As is customary for TES X-ray micro-calorimeter setups, to achieve the optimal spectral performance from the system, the behavior of the detectors under an external magnetic field has to be characterized. This is because the TES critical current is magnetic field dependent, as the system formed by the TES bilayer and the niobium leads constitutes an SS’S Josephson junction. It is typically observed that the best operation is obtained when no magnetic fields perpendicular to the bilayer are present.
These fields might be caused by external sources such as pulse tubes, Earth magnetic field, EMI from laboratory equipment, etc. Furthermore, TESs under dc-bias suffer from the effects of self-field, the magnetic field generated by the loop made by the leads and the TES itself [18]. Such magnetic field adds to the residual field at the TES array and is bias point dependent. The external magnetic fields are typically compensated with a superconducting coil mounted just below the TES array, as also envisaged for X-IFU.
@@
We found the average of the cancelling fields measured with both methods to be consistent. We measured an average cancelling magnetic field at a level of 150 nT with no evident gradient across the array, with a few TESs showing a variation in residual field with respect to neighbouring pixels (see Figure 4). The reason for this variation is not identified, but it could be due to either uncertainty in the measurement or a real effect in the TES array due to non-perfect uniformity. For reference, this value is nevertheless within the T requirement for X-IFU for magnetic-field gradient across the TES array. Furthermore, this value of average residual field is comparable with that typically measured for other TES setups read out under FDM hosted in the same cryostat [19]. The minimization of magnetic field gradients is important for flight operation, because TES pixels within a TDM readout column will share their dc-voltage bias and thus cannot be tuned individually to compensate for eventual different magnetic-field offsets.